OpenStudy (jhannybean):

\[\huge \lim_{x \rightarrow \infty} x^{\frac{3}{\ln(5x^3)}} \]

OpenStudy (alexandervonhumboldt2):

@perl @satellite73

OpenStudy (anonymous):

looks like \(\infty^0\) take the log

OpenStudy (anonymous):

\[\frac{3}{3\ln(5x)}\ln(x)\] is a start

OpenStudy (perl):

$$\Large { \lim_{x \rightarrow \infty} x^{\frac{3}{\ln(5x^3)}} = \lim_{x \rightarrow \infty} e^{\ln x^{\frac{3}{\ln(5x^3)}} } = \lim_{x \rightarrow \infty} e^{\frac{3}{\ln(5x^3)}\cdot \ln x } } $$

OpenStudy (anonymous):

cancel the 3's

OpenStudy (anonymous):

oh no i make a mistake oops

OpenStudy (xapproachesinfinity):

just gonna say that lol

OpenStudy (anonymous):

well we can still break it apart, not sure if it helps or hurts

OpenStudy (jhannybean):

The 3's will cancel there

OpenStudy (anonymous):

\[\frac{3\ln(x)}{\ln(5x^3)}\] is the limit you need to compute

OpenStudy (perl):

$$\Large { \lim_{x \rightarrow \infty} x^{\frac{3}{\ln(5x^3)}} = \lim_{x \rightarrow \infty} e^{\ln x^{\frac{3}{\ln(5x^3)}} } \\ = \lim_{x \rightarrow \infty} e^{\frac{3}{\ln(5x^3)}\cdot ~\ln (x) } = e^{\lim_{x \rightarrow \infty} \frac{3\ln(x)}{\ln(5x^3)}} \\~\\ \Rightarrow^{L'Hopitals} } $$

OpenStudy (anonymous):

you get one pretty much in your head

OpenStudy (perl):

thats an \( \bf \frac{\infty}{\infty} \) indeterminate form

OpenStudy (anonymous):

now i see it is better to bring the exponent in than the other way around \[\frac{\ln(x^3)}{\ln(5x^3)}\]

OpenStudy (jhannybean):

What happened from \[\huge = \lim_{x \rightarrow \infty} e^{\ln x^{\frac{3}{\ln(5x^3)}} } \] to \[\huge = \lim_{x \rightarrow \infty} e^{\frac{3}{\ln(5x^3)}\cdot ~\ln (x) } \]

OpenStudy (perl):

you can save a step and use property of logs

OpenStudy (anonymous):

algebra

OpenStudy (anonymous):

multiply means multiply in the numerator

OpenStudy (rational):

if you wanto work it w/o Lhopital \(u = \ln(5x^3) \implies x = (\frac{e^u}{5})^{1/3}\) the limit becomes \[\lim\limits_{u\to \infty} \left((\frac{e^u}{5})^{1/3}\right)^{3/u} = \lim\limits_{u\to \infty} \frac{e}{5^u} = e \]

OpenStudy (perl):

$$ \Huge e^{\ln(x^a)} = e^{a\cdot \ln(x)} $$

OpenStudy (anonymous):

the exponent is \[\frac{3\ln(x)}{\ln(5x^3)}=\frac{\ln(x^3)}{\ln(5x^3)}\]

OpenStudy (jhannybean):

Ahhh that part is clear now.

OpenStudy (misty1212):

you do not really need l'hopital to compute that limit, but it would work i like @rational method without doing much at all

OpenStudy (jhannybean):

I posted this problem to better understand exp rules, they confuse me a bit.

OpenStudy (perl):

$$ \Large { \lim_{x \rightarrow \infty} x^{\frac{3}{\ln(5x^3)}} = \lim_{x \rightarrow \infty} e^{\ln x^{\frac{3}{\ln(5x^3)}} } \\ = \lim_{x \rightarrow \infty} e^{\frac{3}{\ln(5x^3)}\cdot ~\ln (x) } = e^{\lim_{x \rightarrow \infty} \frac{3\ln(x)}{\ln(5x^3)}} \\\LARGE = e^{\lim_{x \rightarrow \infty} \frac{3\ln(x)}{\ln(5) + 3\cdot \ln (x))}} \\~\\ \Rightarrow^{L'Hopitals} \\\LARGE = e^{\lim_{x \rightarrow \infty} \frac{3 \cdot \frac1x}{0 + 3\cdot \frac1x}} \\\LARGE = e^{\lim_{x \rightarrow \infty} 1} \\\LARGE = e^{1} } $$

OpenStudy (alexandervonhumboldt2):

@rational 's method is shorter but both methods are good.

OpenStudy (jhannybean):

Thanks @perl and @rational :) Cleared things up alot more.

OpenStudy (perl):

some parentheses would make it clearer

OpenStudy (perl):

maybe :)

OpenStudy (jhannybean):

\(\color{blue}{\text{Originally Posted by}}\) @rational if you wanto work it w/o Lhopital \(u = \ln(5x^3) \implies x = (\frac{e^u}{5})^{1/3}\) the limit becomes \[\lim\limits_{u\to \infty} \left((\frac{e^u}{5})^{1/3}\right)^{3/u} = \lim\limits_{u\to \infty} \frac{e}{5^u} = e \] \(\color{blue}{\text{End of Quote}}\) Didn't catch the u-sub that quickly, haha.

OpenStudy (rational):

personally i feel the best way to understand exponents is to study them with logarithms

OpenStudy (xapproachesinfinity):

just a correction to rational's version it is supposed to be 5^(1/u) no 5^u

OpenStudy (perl):

theres a typo there

OpenStudy (perl):

lim e / 5^{1/u}

OpenStudy (xapproachesinfinity):

yes! that's my point :)

OpenStudy (rational):

Ahh right it has to be 1/u

OpenStudy (xapproachesinfinity):

good substitution :)

OpenStudy (perl):

yes thats pretty nifty

OpenStudy (perl):

@rational Is there a way to find limit lim x^x as x ->0+ without Lhopital's rule , analytically

OpenStudy (rational):

that looks interesting you're doing e^(x lnx) hoping to use Lhopital right

OpenStudy (perl):

yeah, i was wondering if there was a clever substitution to avoid Lhopitals rule

OpenStudy (perl):

thanks inky

OpenStudy (inkyvoyd):

yay stack exchange :D

OpenStudy (perl):

ill just finish this problem

OpenStudy (perl):

i need quote addon for mozilla firefox

OpenStudy (xapproachesinfinity):

that's an interesting way to do it it required using some linear approximation

OpenStudy (perl):

$$ \LARGE{ \lim_{x \to 0^+}x^x = \lim_{x \to 0^+}e^{\ln(x^x)} =\\ \lim_{x \to 0^+}e^ { x \cdot \ln x } = \lim_{x \to 0^+} e^ { \frac{\ln x }{ 1/x} } \\ \Rightarrow^{L'Hopital's} \lim_{x \to 0^+} e^ { \frac{\frac1x }{ -1/x^2} } \\=\lim_{x \to 0^+} e^ { \frac 1 x \cdot \frac{-x^2}{1} } \\=\lim_{x \to 0^+} e^ { -x } = e^{-0} =e^{0}= 1 } $$

OpenStudy (perl):

delete what (i was busy with Latex)

OpenStudy (rational):

i liked \(x = e^{-t}\) substittuion reply in that mse link

OpenStudy (rational):

ofcourse that "requires" lhopital to conclude: \[\lim\limits_{t\to\infty} \dfrac{t}{e^t} = 0\] he is just masking it by saying that limit is intuitive enough for everybody

OpenStudy (xapproachesinfinity):

didn't the replay compare that with lim 2/t which we know goes to zero there isn't l'hopital rule use ?

OpenStudy (xapproachesinfinity):

reply*

OpenStudy (rational):

\[e^t \ge t^2/2\] part ?

OpenStudy (xapproachesinfinity):

that linear approximation

OpenStudy (rational):

yes thats nice and clever but still don't you think we can always say \[e^t \gt t^n\] for all \(n\ge 0\) and large enough \(t\) ?

OpenStudy (perl):

$$\Large{ e^t = 1 + t + t^2/2 + ... \\ \therefore \\ e^t \geq t^2/2 } $$

OpenStudy (rational):

i think derivatives/linear approximations/lhopital all are same

OpenStudy (xapproachesinfinity):

yes true, but we needed t^2/2 to ur case so we just ignore other facts lol

OpenStudy (rational):

for limit w/o lhopital, i think we're not allowed to use any of them... dont even think of taylor

OpenStudy (perl):

what do you mean by linear approximation?

OpenStudy (xapproachesinfinity):

hmmm you do have a point! there the use of many theorem that fall into l'hopital

OpenStudy (perl):

i mean how would we use it here

OpenStudy (xapproachesinfinity):

l'hopital if recall came by mean value theorem

OpenStudy (rational):

http://i.stack.imgur.com/GU8wd.jpg lmao

OpenStudy (jhannybean):

heading off for the night, thanks again you guys.

OpenStudy (rational):

im pretty sure its very hard to make sense of the flaw in that fakeproof w/o analysis tools (analysis = rigor)

OpenStudy (perl):

true